### v1

# Improving Relational Database Interactions with Large Language Models: Column Descriptions and Their Impact on Text-to-SQL Performance

Abstract
Relational databases often suffer from uninformative descriptors of table contents, such as ambiguous columns and hard-to-interpret values, impacting both human users and Text-to-SQL models. This paper explores the use of large language models (LLMs) to generate informative column descriptions as a semantic layer for relational databases. Using the BIRD-Bench development set, we created ColSQL, a dataset with gold-standard column descriptions generated and refined by LLMs and human annotators. We evaluated several instruction-tuned models, finding that GPT-4o and Command R+ excelled in generating high-quality descriptions. Additionally, we applied an LLM-as-a-judge to evaluate model performance. Although this method does not align well with human evaluations, we included it to explore its potential and to identify areas for improvement. More work is needed to improve the reliability of automatic evaluations for this task. We also find that detailed column descriptions significantly improve Text-to-SQL execution accuracy, especially when columns are uninformative. This study establishes LLMs as effective tools for generating detailed metadata, enhancing the usability of relational databases.
1 Introduction
Figure 1: Example of a question and golden SQL query pair from the BIRD-Bench text-to-SQL benchmark illustrating the need for external knowledge. For a text-to-SQL system to produce a correct SQL query, it needs to know that column A2 and A3 correspond to the district name and region name respectively, which is not intuitively apparent directly from the original column names.
Structured data in relational databases is crucial for modern information systems. It not only stores essential business data but also allows organizations to analyze trends, optimize operations, and make informed decisions, maintaining a competitive edge in today’s fast-paced economy. In this setting, text-to-SQL, which converts natural language queries into SQL commands, enhances access to structured data by reducing the need for expert knowledge.
However, working with these databases often presents challenges due to complex table schemas and vague descriptors used for table contents, such as ambiguous column names and hard-to-interpret values Talaei et al. (2024). These ambiguities can significantly hinder the effectiveness of both human users and generative models employed in text-to-SQL tasks. A good illustration of such an ambiguity can be found in the district table in the financial database of the BIRD-Bench text-to-SQL benchmark Li et al. (2023), as illustrated in Figure 1. The meaning of columns labeled A2 and A3, which represent the district name and its respective region, are not immediately clear from the names alone.
In response to this challenge, the introduction of descriptive column descriptions as a semantic layer over the existing database structures can dramatically improve their usability. However, manually crafting such column descriptions is a time-consuming and labor-intensive process. The large amount of relational databases in use across society today, underscores the need for automated systems that can generate natural language descriptions for SQL databases in order to enhance their interpretability for both human users and text-to-SQL models.
The recent advancements in LLMs have shown that they can produce coherent and contextually relevant text across various domains, making them potential tools for automating such metadata creation [ref]. By leveraging LLMs, the hypothesis is that the labor-intensive process of generating metadata for relational databases can be significantly simplified. Previous research has used generative LLMs to expand abbreviated column names in a SQL setting Zhang et al. (2023), but to the best of our knowledge, no works exist for generating column descriptions in the general case across a wider array of column types and formats.
In this paper we will for this reason explore how well LLMs can generate column descriptions that contain detailed and relevant information. To enable this study, we first create a dataset for column description generation. We use the development set of BIRD-Bench Li et al. (2023) and its databases as the starting point for our dataset. While BIRD-Bench contains some column descriptions, it does not have complete coverage over all columns, and the existing descriptions are often uninformative.
We first improve the existing column descriptions in BIRD-Bench using an LLM to generate more informative descriptions, demonstrating the usefulness of an LLM in a scenario where some initial metadata exists. Secondly, expert human annotators manually correct the generated descriptions to create a refined dataset of gold-standard references as well as annotating the difficulty of generating a column description for each specific column. This corrected dataset serves as a benchmark for evaluating models on column description generation.
We then apply several instruction-tuned models of different sizes and types to our dataset and perform an extensive human evaluation for all models. We find that both GPT-4o and Command R+ frequently generate column descriptions that have perfect overlap with the gold standard.
In addition to human evaluations, we use an LLM-as-a-judge to evaluate model performance, studying how this automatic method correlates with human annotations and its viability for ranking models on this dataset. We find that the LLM judge does not align well with human evaluations and more work is needed to refine the LLM judge for this task.
To assess the general utility of column descriptions, we apply a model to a text-to-SQL task on the BIRD-Bench development set, both with and without column descriptions. Our results demonstrate that column descriptions aid in more accurately generating SQL queries, underscoring their importance in such tasks, especially for columns identified as having limited information. To further verify the importance of column descriptions in scenarios with limited information, we create an alternative version of the BIRD-Bench development set featuring completely uninformative column names. In this extreme scenario, we find that the inclusion of column descriptions is crucial for text-to-SQL generation.
Main Takeaway: Our study demonstrates that LLMs can be a useful tool for generating column descriptions, and including detailed column descriptions leads to more accurate text-to-SQL query generation, particularly in scenarios where column information is limited or uninformative. Future work will focus on applying this methodology to generate other types of metadata, such as value descriptions, to improve the usability of relational databases for humans and generative LLMs.
Database No. of Tables No. of Columns Easy Medium Hard Very Hard
debit_card_specializing 5 21 17 3 1 0
financial 8 55 21 9 11 13
formula_1 13 94 49 35 10 0
california_schools 3 89 24 32 30 3
card_games 6 115 37 28 50 0
european_football_2 7 199 34 62 103 0
thrombosis_prediction 3 64 56 2 6 0
toxicology 4 11 8 0 2 1
student_club 8 48 39 9 0 0
superhero 10 31 30 1 0 0
codebase_community 8 71 62 4 5 0
Total 75 798 377 (47%) 185 (23%) 218 (27%) 17 (2%)
Table 1: Number of tables and columns, and the distribution of column difficulty across the databases in the ColSQL dataset.
2 Related Work
2.1 Table Understanding and Metadata Augmentation Tasks
In recent years, significant effort has been devoted to simplifying the understanding of tabular data. Table-to-text Zhao et al. (2023b, a); Kasner et al. (2023); Yang et al. (2022); Gong et al. (2019) and table question answering Pal et al. (2023); Xie et al. (2022); Herzig et al. (2020) aim at developing models able to understand structured tabular data and natural language questions to perform reasoning and tasks across tables.
More relevant to our work are previous studies focusing on column type detection and column relationship identification for improving the understanding of tables. Column type detection involves classifying columns into predefined semantic categories, such as dates or currencies, to enhance data interpretation Suhara et al. (2022); Zhang et al. (2020). Column relationship identification focuses on determining the structural relationships between columns, such as primary and foreign keys, to understand structured data in tables Wang et al. (2021); Deng et al. (2022). Unlike these approaches, our method of generating column descriptions using large language models aims to produce natural language explanations of column contents, offering a more comprehensive and contextually rich understanding of the data.
While the work of Zhang et al. (2023) focuses on expanding abbreviated column names to improve database schema readability and table understanding tasks using LLMs, our work goes further by generating full descriptions for column contents and considering a much more diverse range of column names.
2.2 Metadata Usage in Text-to-SQL
The shift from traditional query languages to natural language interfaces through text-to-SQL has been significantly advanced by current LLMs. However, the inherent ambiguities and missing information in database schemas necessitate the incorporation of external knowledge and metadata to improve conversion accuracy Zhang et al. (2024). Previous studies have explored various approaches to address this challenge. Talaei et al. (2024) propose a pipeline that retrieves relevant data and context from a knowledge store to generate accurate SQL queries, leveraging a data store based on the original BIRD-Bench metadata files. Similarly, Dou et al. (2022) enhance the text-to-SQL process by using external domain-specific knowledge, such as mathematical formulas and publicly available resources.
Sun et al. (2024) introduce question-specific database content to improve tuning performance, while Hong et al. (2024) utilize a tailored data expert LLM to provide useful knowledge to the SQL-generating LLM. Petrovski et al. Petrovski et al. (2018) take a different approach by vectorizing entire data columns into embeddings, thus eliminating the need for traditional descriptions and metadata in the text-to-SQL process. These studies collectively validate that incorporating external knowledge and metadata can significantly enhance text-to-SQL systems.
In contrast to these approaches, our work focuses on generating comprehensive SQL column descriptions directly using LLMs. While prior research emphasizes leveraging external knowledge and complex data retrieval processes, we aim to enhance the interpretability and usability of SQL queries by creating detailed descriptions for SQL columns.
3 Column Description Dataset Creation
To enable our study on generating informative column descriptions, we create ColSQL, a new dataset based on the development set of the widely-used text-to-SQL benchmark, BIRD-Bench Li et al. (2023) [^1] BIRD-Bench is licensed under $C$ BY-SA 4.0, permitting transformation and reuse of the dataset.. This dataset consists of 799 columns across 11 different databases and domains, which are presented in Table 1. We selected BIRD-Bench as the basis for our dataset because it allows us to test column description generation across a broad set of domains. Additionally, the natural language queries in BIRD-Bench were designed so that external knowledge (e.g., metadata) is necessary to perform well on the task. In contrast, the WikiSQL Zhong et al. (2017) and SPIDER Yu et al. (2018) text-to-SQL benchmarks primarily uses exact column names in its natural language queries, making them unsuitable for our use case.
While BIRD-Bench already includes column descriptions, we discovered that many columns either lacked descriptions entirely or contained insufficient or erroneous information. This necessitated improving the existing column descriptions to create an adequate gold standard for evaluating models.
In the following sections, we first define what constitutes a good column description. We then describe how we improved BIRD-Bench’s descriptions using generative LLMs and human annotators. Additionally, we explain how we annotated the difficulty level for generating a description from each column.
3.1 What is a good column description?
Natural language descriptions are crucial for addressing the often subtle ambiguities and complexities found in column names. Previous works that relied on a single word or type to describe columns are insufficient for capturing the full context and meaning. A more detailed description ensures that the interpretation of the column is free of ambiguity, facilitating better understanding and more accurate querying.
Defining exactly what constitutes a good description in a set of rules is a complex task, because of the complex nature of natural language and the complexities in SQL databases. However, we used the following criteria to define a good description:
  • It should eliminate ambiguity and enhance the interpretability of a column.
  • It should provide clear and precise information about the column, specifying the nature of the data it holds and its context within a table.
  • It should include information about the subject of the reference, e.g., instead of "The name," use "The name of the client that made the transaction" to ensure clarity.
  • It must be a complete and grammatically correct sentence that explicitly describes what the column represents.
We do not require the inclusion of interpretations of specific values inside the database, leaving the interpretation and generation of metadata for database values for future work.
Given the numerous ambiguities and complexities in SQL databases due to omitted information, we invite the community to engage in discussions about what constitutes a good column description and the necessary metadata to include.
3.2 Improving Existing Column Descriptions with LLMs
Label Original LLM Updated
Perfect 40.24% 70.80%
Somewhat correct 38.80% 18.80%
Incorrect 1.88% 0.06%
I can’t tell 0.81% 7.27%
No description 18.27% 3.07%
Table 2: Quality of the original BIRD-Bench column descriptions and the LLM improved descriptions.
We aimed to improve existing column descriptions using LLMs for two reasons: as a preliminary step before human annotators create the gold standard dataset, and to evaluate the effectiveness of LLMs in improving existing column descriptions.
To improve column descriptions, we used GPT-4o, providing it with the BIRD-Bench column descriptions, corresponding schema in CREATE_TABLE statements, example data rows from the corresponding table, and instructions to generate descriptions for given columns. The model was given clear guidelines and in-context examples of the desired description format. The full prompt template is in Appendix B.4.
Two human annotators who are authors of this paper, fluent in English, and experts in SQL, assessed the original BIRD-Bench descriptions and the LLM-generated descriptions, independently assigning each description one of the following labels:
  • Perfect: Contains enough information so that the interpretation of the column is completely free of ambiguity.
  • Somewhat Correct: Matching the perfect description but verbose with redundant information (although without any incorrect or misleading information).
  • Incorrect: Contains incorrect information.
  • No Description: Description is missing.
  • I can’t tell: Annotator cannot determine the description’s accuracy based on available data.
The full annotator guidelines and more detailed descriptions of the labels are presented in Apendix B.1.2.
Annotators’ assessments of the two datasets are shown in Table 2. There was a substantial level of agreement between annotators, with a Cohen’s kappa of $0.680.68$ on the original BIRD-Bench descriptions and $0.610.61$ on the improved generated descriptions. The GPT-4o descriptions showed a significant improvement over the original BIRD-Bench descriptions. While not perfect, LLM-generated descriptions can be a useful tool for improving existing column metadata when there are already quality issues in the existing column metadata.
Figure 2: Mean quality scores for the generated column descriptions from each model on the ColSQL dataset, as rated by the human evaluators.
Difficulty GPT-4o Command R+ Mixtral 22x8B Qwen 2 72B CodeStral 22B Mistral 7B
Easy 3.83 3.55 3.26 2.98 2.97 3.24
Medium 3.66 3.28 3.18 2.89 2.79 2.84
Hard 3.46 2.46 2.90 2.76 2.69 2.04
Very Hard 1.93 1.24 1.59 1.47 1.44 1.12
Table 3: Mean quality scores for each model’s generated descriptions across the column difficulty levels in the ColSQL dataset, as rated by the human evaluators.
3.3 Creating the Gold Standard Dataset
Since not all LLM-generated improvements were perfect, manual corrections of the generations were necessary to achieve the high standard required for a gold standard dataset. First, disagreements in scores were addressed by reviewing each other’s annotations and then through discussion to reach a consensus. Descriptions that did not achieve a “Perfect” quality score from both annotators after this phase were manually corrected until both annotators agreed on a “Perfect” rating. The final set of descriptions was then reviewed to ensure they provided clear, comprehensive, and contextually rich information for each column. Through this process, we believe we have created a dataset of high quality, serving as a reliable reference for evaluating the performance of models in generating column descriptions. The final composition of the dataset can be found in Table 1. Examples of gold descriptions can be found in Appendix B.3.
3.4 Annotating Column Difficulty
Following the finalization of the gold standard descriptions, each column was assessed on the difficulty in generating descriptions for it. Columns were categorized into four levels: "Easy," "Medium," "Hard," and "Very Hard." These levels reflect the annotators’ assessment of the amount of available information necessary to generate an accurate description for a column. Examples of columns of each difficulty can be found in B.3.
At the "Easy" level, annotators could determine the column description by only looking at the schema of the relevant table. For example, in the Financial database, the Client table’s birth_date column is self-explanatory.
At the "Medium" level, annotators could accurately determine the column description given example data in addition to the table schema from the database. This suggests that while some additional context is required, the column’s role can still be deduced with a reasonable amount of information. For instance, in the student_club database, the zip_code table contains a column named short_state. Given the data inside the database, it is possible to accurately determine that this column contains two-letter abbreviations of states.
Columns categorized as "Hard" present greater challenges. Even with the schema and example data from the database, annotators were unsure about what the column description should be. This uncertainty arises from ambiguities or complexities that cannot be resolved with the given information, highlighting the need for further context or domain knowledge. For example, in the formula_1 database’s results table, the rank column has data indicating rankings, but without context, it is hard to determine if it refers to race position, qualifying rank, or another metric. The description, "The overall rank of the driver’s fastest lap time in the race," clarifies its purpose.
The "Very Hard" level involved columns where accurate descriptions were impossible without additional documentation. An example is the A11 column in the Districts table of the Financial database, which contains only numbers and a completely uninformative name.
The assessment was performed by two independent annotators, distinct from those who created the gold standard descriptions, but with the same level of expertise. The inter-annotator agreement for difficulty level was fair, with a Cohen’s kappa of 0.31.
After completing the annotation process, the annotators used the same procedure as for the gold descriptions to resolve disagreements. It became clear during their discussion that one annotator had misunderstood the difference between the "Easy" and "Medium" criteria, causing the low agreement score. This annotator incorrectly thought "Easy" included only columns with unambiguous names, whereas our definition allows considering other column names in the table. All remaining disagreements were resolved through discussion, resulting in one final difficulty score for each column.
The distribution of columns across the four difficulty levels is presented in Table 1. We observed that a significant portion of columns fell into the “Easy” and “Medium” categories, with “Very Hard” columns primarily residing in the Financial domain, where tables often have completely uninformative column names such as “A1”.
We believe that the difficulty ratings of columns will aid in understanding and improving the performance of models in generating column descriptions across varying levels of complexity.
4 Generating Column Descriptions
In this section, we assess LLMs’ ability to generate detailed column descriptions using our dataset, using both human evaluations and the LLM-as-a-judge.
Annotator/LLM Command R+ CodeStral 22B Mistral 7b Qwen 2 72B Mixtral 22B GPT-4o
Annotator 1 vs GPT-4o 0.22 0.16 0.24 0.21 0.16 0.12
Annotator 2 vs GPT-4o 0.25 0.18 0.23 0.22 0.17 0.12
Annotator 1 vs Annotator 2 0.60 0.71 0.71 0.62 0.63 0.64
Table 4: Cohen’s Kappa scores between annotators, and between annotators and a GPT-4o LLM judge.
4.1 Models
To evaluate the effectiveness of LLMs in generating column descriptions, we used our dataset presented in Section $3.$ We evaluated the following models: GPT-4o, Mixtral-8x22B-Instruct-v0.1 Jiang et al. (2024), Command R+, Qwen2-72B-Instruct Yang et al. (2024), Codestral-22B-v0.1, and Mistral-7B-Instruct-v0.3 Jiang et al. (2023). These models were selected to provide a diverse range of perspectives, including a state-of-the-art closed-source model (GPT-4o), several newly released high-performing open-source models (Qwen2-72B, Mixtral-8x22B, Codestral-22B, Command R+), and a smaller model (Mistral-7B).
Each model was given the database schema, the specified column, example data rows from the corresponding table, and specific instructions for generating the column description. The exact prompt is presented in Appendix B.4.
Inference for the open-source models was performed on several A100 GPUs, using BitsandBytes Dettmers et al. (2022) to load the models in 8-bit format. We set the model temperature to 0.7.
4.2 Human Evaluation
For each model’s generated output, two human annotators independently assessed the quality of the column descriptions. As previously, the set of annotators was different from the ones that created the gold dataset, but with the same type of expertise. The annotations were also based on a slightly different set of criteria from what was used when rating the original and improved column descriptions, presented in Section 3.2, as we now have the gold column descriptions as a reference point.
The generated outputs were scored for their usefulness, using labels “Perfect”, “Almost Perfect”, “Somewhat correct”, “Incorrect”, and “No Description”. For example, the main difference between “Perfect” and “Almost Perfect”, is that the latter contains additional redundant information that is not incorrect. A “Somewhat correct” column description is correct, but there is information missing. The full instructions and definitions of the labels are presented in Appendix B.1.4.
To ensure the reliability of our evaluations, we measured the inter-annotator agreement for each model’s output. Cohen’s kappa for each model is presented in Table 4. The agreement among annotators was above 0.60 for all models, indicating a substantial agreement between the annotators.
The results of the model evaluations are presented in Figure 2 and Table 3. Our observations show that GPT-4o and Command R+ generated the highest proportion of "Perfect" descriptions, indicating their ability to provide concise and accurate column descriptions. In contrast, Qwen 2 72B and Codestral 22B had a high number of "Almost perfect" descriptions, suggesting a tendency to be verbose and add unnecessary information. Incorrect generations were relatively uncommon for most models, which might be more important than the distinctions between different levels of correctness. However, the ratio of incorrect answers were higher for smaller models like Mistral 7B.
4.3 Evaluating Models Using LLM-as-a-Judge
We also performed an evaluation using LLM-as-a-judge, with GPT-4o as the judge. This evaluation aimed to investigate whether there exists an automated metric that could effectively replace human evaluations.
The LLM judge was provided with a task description, the model-generated descriptions, the gold standard references, and the rating criteria. The full prompt is presented in Appendix B.4.
We compared the model rankings as judged by GPT-4o to those determined by human annotators to see if there was a consistent correlation. The results, as shown in Table 4, indicate low alignment between GPT-4o and human annotators, with Cohen’s Kappa scores ranging from 0.12 to 0.25. Notably, the agreement between human annotators was significantly higher, with Kappa scores ranging from 0.60 to 0.71.
As a small side study, we also combined the “Perfect” and “Almost” Perfect’ scores, resulting in a doubling of Kappa values. Although the scores remain modest, this suggests potential for refining the evaluation process to better align with human ratings.
5 Impact of Column Descriptions for Text-to-SQL
To study the usefulness of column descriptions for text-to-SQL systems, we evaluated the best-performing model (GPT-4o) on the development set of BIRD-Bench, which contains 1534 natural language questions and SQL query pairs. Our goal was to isolate and understand the effects of column descriptions on model performance.
The experimental setup involved four scenarios:
  1. 1.
    Without any column descriptions: The model operates with only the database schema.
  2. 2.
    With BIRD-Bench original column descriptions: The model uses the original descriptions provided by BIRD-Bench in addition to the schema.
  3. 3.
    With generated column descriptions: The model uses descriptions generated by the best model from Section 4 in addition to the schema, simulating a realistic scenario.
  4. 4.
    With gold standard descriptions: The model uses the manually refined gold standard descriptions in addition to the schema, representing the best-case scenario.
Included Metadata Execution Accuracy
Without col. descr. 0.30
With BIRD-Bench col. descr. 0.31
With GPT-4o predicted col. descr. 0.32
With gold col. descr. 0.33
Table 5: Execution Accuracy over the BIRD-Bench development set for a zero-shot text-to-sql model using GPT-4o with and without different column descriptions.
In all scenarios, we included only column descriptions as the available metadata alongside the database schema. The full prompt for the text-to-SQL generation is presented in Appendix B.4.
The full results are presented in Table 5. Our findings indicate that column descriptions improve model performance on text-to-SQL tasks. We also observe that the GPT-4o generated column descriptions, which have been generated using only the database schema and example data, is slightly better than the existing BIRD-Bench column descriptions, which were created by human annotators.
We notice an improvement for the model with gold descriptions over the model with access to no column descriptions for questions involving difficult columns, where 34% of uniquely correct text-to-SQL samples referenced columns of a ”Hard“ or ”Very Hard“ difficulty rating. This shows that additional contextual information is crucial. In the next section we perform a more detailed study on how column descriptions can help models correctly predict text-to-SQL questions that reference “Very Hard” columns.
5.1 Analyzing the Effects of Difficult Columns
Figure 3: Execution accuracy with and without gold column descriptions on completely uninformative column names.
To better understand the usefulness of column descriptions in low-information scenarios, we replaced all the column names in each database with uninformative column names. The first table in database A where denoted by A1, the second by A2, and so on.
We chose to study this extreme naming scenario to highlight the crucial role of column descriptions when column names provide little to no context. Studying this extreme case is important because it helps us understand the boundaries of model performance and the essential need for descriptive metadata. While such uninformative column names occur in the BIRD-Bench dataset, most realistic scenarios likely fall somewhere between these uninformative names and more informative column names.
The results of this experiment are presented in Figure 3. The findings clearly show that column descriptions are highly useful when performing text-to-SQL in scenarios with low-information column names, as the model’s performance significantly improves when detailed descriptions are provided.
Through this experiment, we illustrate that detailed column descriptions are essential for the robustness of text-to-SQL systems, especially in challenging scenarios where column names alone are insufficient.
6 Conclusions and Future Work
In this study, we have created ColSQL, a dataset with curated human annotations of column descriptions. This dataset serves as a reliable gold standard for evaluating the performance of models in generating accurate and contextually rich column descriptions.
Our main findings demonstrate that LLMs can effectively generate high-quality column descriptions. Among the models evaluated, GPT-4o and Command R+ performed best, particularly in generating descriptions that meet our "Perfect" criteria where GPT-4o meets the criteria for 75% of the columns.
Through our evaluation process, supported by rigorous human annotations, we have highlighted both the strengths and limitations of various models in this task and shown that LLM-as-a-judge is not yet viable alternative to human annotations when ranking models for this task. We leave it to future work to find reliable automatic evaluations for this task.
We have also shown that detailed column descriptions significantly enhance text-to-SQL systems, especially for columns that are completely uninformative. For uninformative columns we see a 20% increase. This underscores the importance of metadata as a bridge between the database system and generative models interacting with that system.
Future work will explore the generation of other types of metadata, such as table, value, and relationship descriptions, and their impact on text-to-SQL systems. By enhancing metadata, we aim to further improve the accuracy and usability of text-to-SQL systems.
In conclusion, the creation and utilization of detailed metadata is essential for the development of text-to-SQL models and for enabling human users of database systems. This paper is a first step in demonstrating the role of LLMs as generators of such metadata.
Limitations
In this study, we restricted our evaluation to the development set of BIRD-Bench due to the labor costs associated with performing the multitude of human evaluations required. Although the full training set encompasses more domains and might offer a broader evaluation scope, we believe that the development set contains sufficient variety to ensure robust testing of the models. The chosen subset still represents a wide range of domains, allowing us to draw meaningful conclusions about the performance of large language models in generating column descriptions.
Another limitation is that our annotators were not experts in all the specific domains covered by the dataset, such as Thrombosis Prediction. Despite this, the annotators reported confidence in their ability to accurately label column descriptions across all domains, including these more challenging areas. Nevertheless, the lack of domain-specific expertise could introduce some bias or inaccuracies in the annotation process. Future work should consider incorporating domain experts for annotation, particularly for specialized fields, to further enhance the reliability of the dataset.
Ethical Statement
All annotators involved in this study were fully informed about the workload and had the option to decide whether they wanted to participate. Since all annotators were also authors of this paper, they were not compensated for their efforts. However, we ensured that the annotations were produced ethically, with voluntary participation and informed consent from all contributors.
It is important to note that the metadata generation models presented in this paper are not perfect. A model that achieves high scores on our dataset may not generalize well to scenarios outside of the domains we studied. Consequently, caution should be exercised when using these models to generate column descriptions in systems that will be used by humans or other AI systems, such as text-to-SQL models. Ensuring the reliability and accuracy of generated metadata in different contexts is crucial to prevent potential misinterpretations or errors in downstream applications.
Appendix A Datasheet for Datasets
A.1 Motivation
For what purpose was the dataset created?
Relational databases often suffer from uninformative descriptors of table contents, such as columns and values, impacting both human users and text-to-SQL models. This dataset was created to explore and assess how large language models can generate detailed and meaningful column descriptions to be used as a semantic layer over databases, to enhance their usability.
Who created this dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
This question will be answered after the anonymity period.
What support was needed to make this dataset?
This question will be answered after the anonymity period.
A.2 Composition
What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)?
The dataset is made up of 11 SQLite databases, accompanied by a csv file which contains gold standard column descriptions and difficulty ratings for every column in the entire dataset. Each instance therefore represents a column and its accompanying gold description and difficulty rating.
How many instances are there in total (of each type, if appropriate)?
There are 798 columns in total across the 11 databases in the dataset. A full breakdown of columns over the different domains can be found in Table 1.
Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?
The dataset is based on the BIRD-Bench development set and includes all databases, tables, and columns from the development set.
What data does each instance consist of?
Each instance contains a column, its corresponding table and database, and its accompanying gold description and difficulty rating.
Is there a label or target associated with each instance? If so, please provide a description.
Yes. Each column contains a corresponding gold column description, as well as an annotated difficulty rating (Easy, Medium, Hard or Very Hard). These levels reflect the annotators’ assessment of the amount of available information necessary to generate an accurate description for a column. Examples of columns of each difficulty can be found in B.3.
Is any information missing from individual instances?
No. The dataset has full coverage over the entire set of columns from the included databases.
Are there any errors, sources of noise, or redundancies in the dataset?
The dataset was created using a combination of automated techniques and human annotators. Although the goal was to create a dataset free of noise, it could be that some noise exists inadvertently due to human error.
Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?
The dataset is self-contained.
Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor-patient confidentiality, data that includes the content of individuals’ non-public communications)?
No.
Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?
No.
A.3 Collection
How was the data associated with each instance acquired? Was the data directly observable (e.g., raw text, movie ratings), reported by subjects (e.g., survey responses), or indirectly inferred/derived from other data (e.g., part-of-speech tags, model-based guesses for age or language)? If data was reported by subjects or indirectly inferred/derived from other data, was the data validated/verified? If so, please describe how.
The data was directly observable and derived from the BIRD-Bench development set. The column descriptions were generated using a large language model and then manually verified and corrected by human annotators. The difficulty ratings where annotated and verified by human annotators.
Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created. Finally, list when the dataset was first published.
The data was collected and the column descriptions were generated and annotated over a period of three months. The dataset will be published after the anonymity period.
What mechanisms or procedures were used to collect the data (e.g., hardware apparatus or sensor, manual human curation, software program, software API)? How were these mechanisms or procedures validated?
The dataset was collected from the BIRD-Bench development set. LLMs were used to generate initial column descriptions, which were then manually curated and corrected by human annotators. The validation was performed by expert annotators through a consensus process.
What was the resource cost of collecting the data? (e.g., what were the required computational resources, and the associated financial costs, and energy consumption - estimate the carbon footprint.
The resource cost included the computational expense of running LLMs on multiple A100 GPUs and the human labor for manual annotation. The exact financial cost and carbon footprint would require further detailed calculations, but they are comparable to typical machine learning model training and annotation tasks.
If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
The dataset is not a sample; it includes the complete BIRD-Bench development set.
Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
The data collection and annotation were performed by the authors of the paper, who are researchers with expertise in SQL and natural language processing. There was no additional compensation as the work was part of their research activities.
Were any ethical review processes conducted (e.g., by an institutional review board)? If so, please provide a description of these review processes, including the outcomes, as well as a link or other access point to any supporting documentation.
No formal ethical review was conducted as the data does not involve personal or sensitive information.
Does the dataset relate to people? If not, you may skip the remainder of the questions in this section.
No.
A.4 Preprocessing / Cleaning / Labeling
Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)? If so, please provide a description. If not, you may skip the remainder of the questions in this section.
Yes, preprocessing included several steps:
  • Data cleaning steps were applied to the original BIRD-Bench databases and dataset:
    • Removed non "utf-8" tokens from the original BIRD-bench metadata CSV files.
    • Corrected spelling in the european_football_2 database in the "Country" description header from "desription" to "description".
    • Removed all columns with zero data and the name "Unnamed" due to (too many/missing) commas in the original BIRD-Bench metadata CSV files.
    • Changed the file name of "ruling.csv" to "rulings.csv" to match the original table name in the card_games domain.
    • Changed "set_transactions.csv" to "set_translations.csv" to match the database name in the card_games domain.
    • Removed all .DS_STORE files from the data directory.
    • Changed the names of the CSV files in the student_club database to match the original table names. Changed upper case to lower case on the first letter of all CSV files, and corrected the code in Zip_Code.
    • Removed the "wins" column from constructors.csv file as the column does not exist in the formula_1 database.
  • Generating initial column descriptions using LLMs, followed by manual correction and labeling of the descriptions.
  • Categorizing the columns based on the difficulty of generating their descriptions.
Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the “raw” data.
Yes, the raw BIRD-Bench development set data was saved and is available upon request.
Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point.
Yes, the code used for preprocessing, cleaning, and labeling is available at https://github.com/anonymous (to be updated post anonymity period).
A.5 Uses
Has the dataset been used for any tasks already? If so, please provide a description.
The dataset has been used to evaluate the performance of various large language models in generating column descriptions and to study the impact of these descriptions on text-to-SQL systems.
Is there a repository that links to any or all papers or systems that use the dataset? If so, please provide a link or other access point.
This will be updated after the anonymity period with the appropriate links.
What (other) tasks could the dataset be used for?
The dataset could be used for tasks such as evaluating and improving text-to-SQL models, studying the impact of metadata on database usability, and developing automated systems for generating database documentation.
Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses? For example, is there anything that a future user might need to know to avoid uses that could result in unfair treatment of individuals or groups (e.g., stereotyping, quality of service issues) or other undesirable harms (e.g., financial harms, legal risks)? If so, please provide a description. Is there anything a future user could do to mitigate these undesirable harms?
The dataset is composed of data generated and annotated for research purposes and does not include personal or sensitive information. Future users should ensure that the generated metadata is appropriate for their specific use case to avoid misinterpretations or errors in applications.
Are there tasks for which the dataset should not be used? If so, please provide a description.
The dataset should not be used for applications involving sensitive or personal data without proper validation and consideration of privacy implications.
A.6 Distribution
Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created? If so, please provide a description.
Yes, the dataset will be made publicly available for research purposes.
How will the dataset be distributed (e.g., tarball on website, API, GitHub)? Does the dataset have a digital object identifier (DOI)?
The dataset will be distributed via GitHub and will be assigned a DOI for reference.
When will the dataset be distributed?
The dataset will be distributed after the anonymity period.
Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)? If so, please describe this license and/or ToU, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms or ToU, as well as any fees associated with these restrictions.
The dataset will be distributed under the Creative Commons Attribution-ShareAlike 4.0 International License ($C$ BY-SA 4.0).
Have any third parties imposed IP-based or other restrictions on the data associated with the instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any relevant licensing terms, as well as any fees associated with these restrictions.
No.
Do any export controls or other regulatory restrictions apply to the dataset or to individual instances? If so, please describe these restrictions, and provide a link or other access point to, or otherwise reproduce, any supporting documentation.
No.
A.7 Maintenance
Who is supporting/hosting/maintaining the dataset?
This will be answered after the anonymity period.
How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
This will be answered after the anonymity period.
Is there an erratum? If so, please provide a link or other access point.
Not at this time.
Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)? If so, please describe how often, by whom, and how updates will be communicated to users (e.g., mailing list, GitHub)?
The dataset will be updated as necessary to correct any errors or add new instances. Updates will be communicated via the GitHub repository.
If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were individuals in question told that their data would be retained for a fixed period of time and then deleted)? If so, please describe these limits and explain how they will be enforced.
Not applicable.
Will older versions of the dataset continue to be supported/hosted/maintained? If so, please describe how. If not, please describe how its obsolescence will be communicated to users.
Older versions of the dataset will continue to be available on GitHub. Any deprecation or updates will be clearly communicated through the repository’s documentation.
If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so? If so, please provide a description. Will these contributions be validated/verified? If so, please describe how. If not, why not? Is there a process for communicating/distributing these contributions to other users? If so, please provide a description.
Contributions can be made through pull requests on the GitHub repository. Contributions will be reviewed and validated by the dataset maintainers before being accepted and merged. All updates and contributions will be documented in the repository’s changelog.
A.8 Any other comments?
None at this time.
Appendix B Appendix
B.1 Annotation Procedure and Guidelines
B.1.1 Annotating the Quality of the Original BIRD-Bench Dataset Column Descriptions
Two annotators independently annotated the column descriptions from the BIRD-Bench dataset, following the predefined annotation guidelines. Before starting the annotation process, an introductory session was held, during which a few samples were annotated collaboratively and discussed with a third party. This session aimed to clarify the decision boundaries within the annotation guidelines. During the annotation process, the annotators had access to the SQLite database and SQLite-viewer, the database name, the table name, the column name, examples of annotations of each category and the metadata files provided by BIRD-Bench. When annotating a column, the annotators were first asked to look at the column name and the table name before reading the column description in the corresponding BIRD-Bench metadata file. After reading the description, the annotators were then asked to also take a look at the corresponding database and table and accompanying data in SQLite Viewer. With this whole picture in mind, the annotators were then instructed to determine the quality of the column description according to the guidelines.
Classification Description
Perfect A perfect column description should contain enough information so that the interpretation of the column is completely free of ambiguity. It does not need to include any descriptions of the specific values inside the column to be considered perfect. The description should contain information about what table the column is referencing. For example, instead of "The name," we want "The name of the client that made the transaction" if we have a transaction database with columns such as NAME, AMOUNT, and DATE to resolve the ambiguity of what the name refers to. Additionally, the column description should be a full and valid English sentence, with proper grammar, capitalization, and punctuation. For instance, instead of "nationality of drivers" when each instance refers to only one driver, it should be "The nationality of a driver."
Somewhat Correct The column description is somewhat correct, but there is room for improvement.
Incorrect The column description is incorrect. Contains inaccurate or misleading information. It could still contain correct information, but any incorrect information automatically leads to an incorrect rating.
No Description The column description is missing.
I Can’t Tell It is impossible to tell the class of the description with the given information.
Language You are allowed to translate textual data in another language in order to determine the difficulty. The act of translating should not affect the assessment of the difficulty.
Tools SQLite Viewer in VSCode and SQLite3 from the command line. Gold column descriptions, database names, table names, and column names from our gold dataset.
Table 6: The annotators’ guidelines for classifying the original column descriptions provided by the BIRD dataset.
B.1.2 Annotating the Quality of the Improved Column Descriptions
The same annotation procedure that was used for annotating the quality of the original BIRD-Bench descriptions was used for annotating the improved columns, which can be found in Section B.1.1.
B.1.3 Annotating the Difficulty of Predicting the Column Descriptions for the BIRD Dataset
This section describes the annotations process used when annotating the difficulty of the columns. During the initial annotation phase, both annotators were given access to the necessary tools, including SQLite Viewer in VSCode and SQLite3 from the command line. Each annotator independently reviewed the database name, table name, column name, example data, gold column description and other columns in the table. Based on this information, they assigned a difficulty level to the task of writing a column description. Annotations were documented in a shared spreadsheet. After the initial annotation phase, the annotators compared their ratings. For columns with matching difficulty ratings, the results were finalized. Discrepancies were discussed, referring to the guidelines and data examples to reach a consensus. If consensus could not be reached, a mediator reviewed the case and made the final decision. The final difficulty ratings were documented in the shared spreadsheet, reflecting the consensus or mediator decisions for all columns creating the final gold standard difficulty annotations.
Difficulty Level Description
Very Hard Given the database name, the table name, the column name and example data from the database, and other columns in the table, it is impossible to accurately determine what the column description should be.
Hard Given the database name, the table name, the column name and example data from the database, and other columns in the table, I am unsure what the column description should be.
Medium Given the database name, the table name, the column name and example data from the database, and other columns in the table, I can accurately determine what the column description should be.
Easy Given only the table name and the column name, and other columns in the table, I can accurately determine what the column description should be.
Language You are allowed to translate textual data in another language in order to determine the difficulty. The act of translating should not affect the assessment of the difficulty.
Tools SQLite Viewer in VSCode and SQLite3 from the command line. Gold column descriptions, database names, table names and column names from our gold dataset.
Table 7: Annotation Guidelines for Difficulty Ratings.
B.1.4 Annotating the Quality of Generated Column Descriptions
In this section, the annotation procedure to evaluate the quality of the descriptions generated by the different LLMs are described. First, four annotators were organized into two teams of two, with each team assigned to three different LLMs. The guidelines categorized the quality of the column descriptions into five levels: Perfect, Almost Perfect, Somewhat correct, Incorrect, and No description. Before starting the annotation process, an introductory session was held, during which a few samples were annotated collaboratively and discussed with a third party. This session aimed to clarify the decision boundaries within the annotation guidelines. During the annotation process, the annotators had access to the column name, corresponding table and database name, the golden column description, the decision tree depicted in Figure 4 and the predicted column description. During the annotation process, each annotator independently reviewed the descriptions generated by the LLMs. Annotations were recorded in a shared spreadsheet. Annotation scores were then calculated to quantify the agreement between the annotators and the accuracy of the LLM-generated descriptions relative to the gold standard.
Figure 4: A decision tree designed to help annotators in deciding the quality of the generated column descriptions.
Quality Level Description
Perfect Matching the gold description without extra, redundant information. Redundant information is categorized as descriptions that do not provide useful additional information. For example, "<Gold description> + ’is a primary/foreign key’" (NOT REDUNDANT) versus "<Gold description> + ’is useful for retrieving data’" (REDUNDANT).
Almost Perfect Matching the gold description but verbose with redundant information, without any incorrect or misleading information.
Somewhat correct The column description is somewhat correct but has room for improvement due to missing information. For example, "The Time column records the specific time at which a transaction occurred, formatted in a 24-hour $H$:$M$ pattern," which lacks enough information to make a valid prediction beyond the primary purpose.
Incorrect The column description is incorrect and contains inaccurate or misleading information. Any incorrect information automatically leads to an incorrect rating, even if some correct information is present.
No description The column description is missing. The models are instructed to output "Not enough information to make a valid prediction" if they cannot generate a suitable column description.
Table 8: Annotation Guidelines for LLM Generated Descriptions
B.2 Annotation Scores
Descriptions Agreement Percentage Cohen’s Kappa
Original Descriptions 0.79 0.68
Improved Descriptions 0.82 0.61
Table 9: Agreement Percentage and Cohen’s Kappa for the quality annotations for both the original BIRD-Bench descriptions and our improved generated descriptions (before corrections).
Difficulty Annotation Agreement Percentage Cohen’s Kappa
Difficulty Rating 0.53 0.31
Table 10: Agreement Percentage and Cohen’s Kappa for Difficulty Ratings
B.3 Examples of Columns, Gold Descriptions and Difficulties
Table 11: Illustrative Examples of Column Descriptions by Difficulty
Difficulty Database Name Table Name Column Name Description Why Ranked This Difficulty
Easy superhero gender gender The name of the gender of the superhero. Self explanatory.
Easy student_club event event_date The date the event took place or is scheduled to take place. Self explanatory.
Medium student_club zip_code short_state Two-letter state abbreviation. While short_state suggests state abbreviations, it requires looking at the data to confirm it represents two-letter state codes, thus needing some context.
Medium formula_1 results laps The number of laps completed by the driver in a race. Requires understanding of the racing domain, intuitive for basic knowledge of sports.
Hard card_games foreign_data uuid Unique identifier referencing the related card. Concept of a UUID may not be clear, and connection to "related card" needs understanding of data relationships.
Hard formula_1 results rank The overall rank of the driver’s fastest lap time in the race. Without context, rank could refer to different types of rankings (e.g., race position, qualifying rank), making it ambiguous without more information.
Very Hard financial district A4 The number of inhabitants in the district. Column name provides no intuitive clue. Data only contains numbers giving no clue as to the nature of the data.
Very Hard financial district A11 Numeric code for the district. The column name A11 in the Districts table is uninformative and, with only numeric data, it’s impossible to discern its purpose without documentation.
B.4 Prompt Templates
1"""
2###TASK:
3Context-GenerateColumnDescriptionforDatabase,togiveusersaneasier
4timeunderstandingwhatdataispresentinthecolumn.
5
6DatabaseSchemaDetails:
7""
8{database_schema}
9""
10
11Hereisexampledatafromthetable{table}:""
12
13{example_data}
14
15Hereisupto10possibleuniquevaluesforthecolumn{column}fromthetable
16{table}:
17
18{unique_data}
19
20""
21
22Thecolumnnamefor{column}is{column_name}.Thisisthenameofthecolumn,
23itcancontainimportantinformationaboutthecolumn,andshouldbeusedto
24writethedescription.
25Thepreviouscolumndescriptionis{column_description}.Thisistheold
26descriptionofthecolumn,thisissometimeslackingandshouldbereadand
27rewritten.
28
29###Task
30Generateaprecisedescriptionforthe{column}columninthe{table}table.
31Yourdescriptionshouldinclude:
32-Primarypurposeofthecolumn.Ifthedetailsintheschemadonotsuffice
33toascertainwhatthedatais,return:"Not enough information to make a
34valid prediction."
35Optionally,yourdescriptioncouldalsoinclude:
36-Additionalusefulinformation(ifapparentfromtheschema),formatted
37asanewsentence,butnevermorethanone.Ifnousefulinformationis
38availableorifthedetailsintheschemadonotsufficetoascertain
39usefuldetails,returnnothing.
40
41###Requirements
42-Focussolelyonconfirmeddetailsfromtheprovidedschema.
43-Keepthedescriptionconciseandfactual.
44-Excludeanyspeculativeoradditionalcommentary.
45-DONOTreturnthephrase"in the {table} table"inyourdescription.
46Thisveryimportant.
47
48**Examples:**
49-Foracolumnnamed"no. of municipalities with inhabitants < 499,"
50thedescriptionshouldbe:"This is the number of municipalities with
51fewer than 499 inhabitants."
52-Foracolumnitisbettertodescribedasshortaspossible:So,
53"Details about the ratio of urban inhabitants."ispreferredover"This
54column provides information on details about ratio of urban inhabitants,".
55-Ifthenameis"Frequency",thedescriptionshouldbe:" The frequency
56of transactions on the account"sincethiscomesfromtheaccounttable.
57-Ifthenameis"amount of money"thedescriptionsshouldbe:"
58The amount of money in the order"sincethiscomesfromtheordertable.
59
60###Pleaseskipthe"data_format"andfocussolelyonupdatingthe
61"column_description".
62
63DONOTreturnanythingelseexceptthegeneratedcolumndescription.
64Thisisveryimportant.Theanswershouldbeonlythegeneratedtext
65aimedatdescribingthecolumn.
66"""
Figure 5: The prompt used to improve upon BIRD-Bench already existing column descriptions.
1"""
2###Context-GenerateColumnDescriptionforDatabase,togiveusersaneasier
3timeunderstandingwhatdataispresentinthecolumn.
4
5DatabaseSchemaDetails:
6""
7{database_schema}
8""
9
10Hereisexampledatafromthetable{table}:""
11
12{example_data}
13
14Hereisupto10possibleuniquevaluesforthecolumn{column}fromthetable
15{table}:
16
17{unique_data}
18
19""
20
21###Task
22Generateaprecisedescriptionforthe{column}columninthe{table}table.
23Yourdescriptionshouldinclude:
24-Primarypurposeofthecolumn.Ifthedetailsintheschemadonotsuffice
25toascertainwhatthedatais,return:"Not enough information to make a
26valid prediction."
27Optionally,yourdescriptioncouldalsoinclude:
28-Additionalusefulinformation(ifapparentfromtheschema),formattedas
29anewsentence,butnevermorethanone.Ifnousefulinformationis
30availableorifthedetailsintheschemadonotsufficetoascertainuseful
31details,returnnothing.
32
33###Requirements
34-Focussolelyonconfirmeddetailsfromtheprovidedschema.
35-Keepthedescriptionconciseandfactual.
36-Excludeanyspeculativeoradditionalcommentary.
37-DONOTreturnthephrase"in the {table} table"inyourdescription.
38Thisveryimportant.
39
40DONOTreturnanythingelseexceptthegeneratedcolumndescription.
41Thisisveryimportant.Theanswershouldbeonlythegeneratedtextaimedat
42describingthecolumn.
43"""
Figure 6: The prompt used for generating columns, given only the schema and sample data from the database.
1"""
2###TASK:
3Context-Youareanexpertinconvertingnaturallanguagequestionsand
4instructionsintotheircorrespondingSQLcounterpart.
5
6Databaseschemaandtheassociatedcolumndescriptionsforeachtable:
7
8{database_schema}
9
10UsingvalidSQL,answerthefollowingquestionbasedonthetablesprovided
11abovebyconvertingthenaturallanguagequestionintothecorrespondingSQL
12query.
13
14Question:{question}
15
16###Requirements
17DONOTreturnanythingelseexcepttheSQLquery.Donotthinkoutloud.
18ONLYreturntheSQLquery,nothingelse.
19"""
Figure 7: Zero-Shot Prompting Template.
1"""
2Youareevaluatingaresponsethathasbeensubmittedforaparticulartask,
3usingaspecificsetofstandards.Belowisthedata:
4[BEGINDATA]
5***
6[Task]:ThetaskistogenerateaccuratedescriptionsofcolumnsinSQL
7databases,givenonlyaccesstotheschemainaCREATE_TABLEformatandexample
8rowsfromthedatabase.
9
10Thegoalistocreateinformativedescriptionswhichreducesambiguityand
11increasesunderstandingforusersofthedatabase.
12
13***
14[Submission]:{response}
15***
16[GoldAnswer]:␣␣{gold_answer}
17***
18[Criterion]:EvaluationCriteria
19
20Correctness:
214:Perfect(MatchingtheGOLDdescriptionorbetter):
22
23Matchingthegolddescriptionwithoutextra,redundant
24information.Toredundantinformation,descriptionswhichdonotprovideuseful
25additionalinformation,iscategorized.
26
27Example:<Golddescription>+"is a primary/foreign key."canbeconsidered
28usefulsotheextrainformationisNOTREDUNDANT.
29
30Golddescription" + "isusefulforretrieveingdata" does not any extra useful
31information so is considered REDUNDANT.
32
333: Almost Perfect
34
35Matching the GOLD description but is verbose with extra, redundant information
36(but all the information is correct, ie it does not contain any incorrect,
37misleading information).
38
392: Somewhat correct
40
41The column description is somewhat correct, but there is room for improvement
42due to missing information.
43
44Example: "TheTimecolumnrecordsthespecifictimeatwhichatransaction
45occurred,formattedina24-hourHH:$M$:SSpattern.Notenoughinformationto
46makeavalidpredictionbeyondtheprimarypurpose."
47
481: Incorrect
49
50The column description is incorrect. Contains innacurrate or misleading
51information. Could still contain correct information but any incorrect
52information automatically leads to an incorrect rating.
53
54***
55[END DATA]
56
57Does the submission meet the criterion? First, write out in a step by step
58manner your reasoning about the criterion to be sure that your conclusion is
59correct. Avoid simply stating the correct answers at the outset.
60
61Your response must be RFC8259 compliant JSON following this schema:
62
63{{"reasoning": str, "correctness": int}}
64
65Make sure your output is a valid json string in the format provided above.
66""""
Figure 8: Template for the LLM-as-a-judge experiments.